Data sets A and B are dependent. Find the test statistic.

Assume that the paired data came from a population that is normally distributed.
Answer #1

39)

Calculate the difference of means by subtracting each respective element from both data s X01 - XO2 = 6.1 - 8.5 = -2.4 x11 -

= d Sa/Vn 1.7329166165745/15 = -0.94 = -0.94 7 1.7329166165745/2.236067977499

test statistic is

1.212928

1.212

option second is correct.

please like ??